# Collectively enhanced ground-state cooling in subwavelength atomic arrays

### v2

@@ Acknowledgements.
O.R.-B. acknowledges support from Fundación Mauricio y Carlota Botton and from Fundació Bancaria “la Caixa” (LCF/BQ/AA18/11680093). R.H. acknowledges funding by the Austrian Science Fund (FWF) 10.55776/W1259. SFY thanks the NSF throughPHY-2207972, the CUA PFC PHY-2317134, and the Q-SEnSE QLCI OMA-2016244. A. A.-G. acknowledges support from the AFOSR through their Young Investigator Prize (No. 21RT0751), from the NSF through their CAREER Award (No. 2047380), and from the David and Lucile Packard Foundation. C.C.R. acknowledges support from the European Union’s Horizon Europe program under the Marie Sklodowska Curie Action LIME (Grant No. 101105916). This research was supported in part by the Austrian Science Fund (FWF) [10.55776/COE1] and by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Free-space dipole-dipole interactions
where $\gamma_{0}=\mu_{0}\omega_{0}^{3}|\boldsymbol{\wp}|^{2}/(3\pi\hbar c)$ is the single-atom spontaneous emission rate in free space, $\hat{\boldsymbol{\wp}}=\boldsymbol{\wp}/|\boldsymbol{\wp}|$ is the unit vector pointing in the direction of the atomic dipole matrix element $\boldsymbol{\wp}$. In $\boldsymbol{G}(\mathbf{r})$ is the free space electromagnetic Green’s tensor. The general expression for the Green’s tensor everywhere except at the origin reads
@@ Appendix B Derivation of the effective master equation
In this Appendix, we derive the effective master equation (16) describing the laser cooling of the motional degrees of freedom of the system. The derivation relies on two approximations. First, we assume the internal or spin degrees of freedom to be only weakly excited, which is valid for weak and off-resonant drives. This allows to approximate the optical response of the atomic ensemble by its linear response. Second, that the coupling between internal and external degrees of freedom is smaller than the characteristic timescale of the spin dynamics, which allows for an adiabatic elimination of the spins. Below, we will derive specific condition for the validity of these two assumptions. Note that the derived effective master equation includes the effects arising from light-induced dipole-dipole interactions mediated by the vacuum electromagnetic field, and therefore describes collective cooling in dense atomic ensembles.
@@ B.1 Displacement transformation in the low spin-excitation regime
The starting point of our analysis is the master equation for the full system in the Lamb-Dicke approximation, Eq. (7). Under the assumption of weak saturation, $\langle\hat{\sigma}_{i}^{\dagger}\hat{\sigma}_{i}\rangle\ll 1$ $\forall i$, we approximate the spin operators by bosonic operators with $[\hat{\sigma}_{i},\hat{\sigma}_{j}^{\dagger}]=\delta_{ij}$. Next, we apply the unitary transformation
[…]
that displaces the bosonic spin and phonon operators as $\hat{\sigma}_{i}\rightarrow\hat{\sigma}_{i}+s_{i}$ and $\hat{b}_{i}\rightarrow\hat{b}_{i}+\beta_{i}$. Equation (47) corresponds to a shift of system’s degrees of freedom to a new equilibrium position determined by the strength of the external driving laser. The displacements can be approximated as a power series in the Lamb-Dicke parameter. To leading order they read
@@
where $\hat{H}_{S}$ is defined in Eq. (10) and the last term represents and additional correction of order $\eta^{2}$ arising from the displacement of the atomic position. Note that even for $\beta_{i}^{(1)}=0$, $\hat{H}_{S}$ contains a second order corrections in $\eta$ that arises from the zero-point motion of the emitters. Formally, this correction is obtained from terms of the form $\hat{b}_{i}\hat{b}_{i}^{\dagger}$, which become $\hat{b}_{i}^{\dagger}\hat{b}_{i}+1$ upon normal ordering [see discussion at the end of Sec. II.1]. Finally, the population recycling term reads
@@
where ±$\mathcal{L}_{\pm}\equiv-(i/\hbar)\left[\hat{H}_{\pm},\hat{\mu}\right]+\mathcal{D}_{\pm}\hat{\mu}$. To first order in $\eta$, the interaction reads
@@
$\mathcal{D}_{\pm}\hat{\mu}$ $=\sum_{i,j\neq i}\gamma_{ij}^{\prime}\left[\eta_{i}\left(\hat{\sigma}_{j}\hat{\mu}\hat{\sigma}_{i}^{\dagger}+s_{j}^{(0)}\hat{\mu}\hat{\sigma}_{i}^{\dagger}+s_{i}^{*(0)}\hat{\sigma}_{j}\hat{\mu}\right)\hat{b}_{i}^{\pm}-\eta_{j}\hat{b}_{j}^{\pm}\left(\hat{\sigma}_{j}\hat{\mu}\hat{\sigma}_{i}^{\dagger}+s_{j}^{(0)}\hat{\mu}\hat{\sigma}_{i}^{\dagger}+s_{i}^{*(0)}\hat{\sigma}_{j}\hat{\mu}\right)\right].$ (59b)
@@ B.2 Adiabatic elimination of spin degrees of freedom
$\eta_{i}\Omega_{i},\eta_{i}J_{ij}^{\prime}|s_{j}^{(0)}|,\eta_{i}\gamma_{ij}^{\prime}|s_{j}^{(0)}|\ll\nu_{i},\gamma_{\lambda}\quad\forall i,\lambda.$ (60)
@@
$\begin{split}\mathcal{L}_{\mathrm{int}}(t)\hat{\mu}=&\left[{\mathcal{L}_{+}(t)+\mathcal{L}_{-}(t)}\right]\hat{\mu}\\ =&\sum_{i}\left({e^{i\nu_{i}t}\mathcal{L}_{+,i}+e^{-i\nu_{i}t}\mathcal{L}_{-,i}}\right)\hat{\mu},\end{split}$ (61)
where we defined ±$\mathcal{L}_{\pm}=\sum_{i}\mathcal{L}_{\pm,i}$, with ±$\mathcal{L}_{\pm,i}$ containing only motional operators of emitter $i$.
To derive an effective equation of motion for the atomic motion, we need to define the relevant subspace where the effective dynamics takes place. Accordingly, we define a superoperators $\mathcal{P}$ that projects the dynamics into the subspace where the internal degrees of freedom are in their displaced ground state |$\ket{0}$, $\mathcal{P}\hat{\mu}=\operatorname{Tr}_{S}[\hat{\mu}]\otimes\left|{0}\middle\rangle\middle\langle{0}\right|\equiv\hat{\rho}\otimes\left|{0}\middle\rangle\middle\langle{0}\right|$. We then apply the projector operator formalism for open quantum systems [72, 37, 38, 73] and derive the Nakajima-Zwanzig equation. Under the time scale separation defined by Eq. (60), we approximate the Nakajima-Zwanzig equation to second order in the interaction Liouvillian (58) as
@@
×$\times\int_{0}^{\infty}d\tau\,e^{\tau(\mathcal{L}_{S}-il\nu_{i})}\mathcal{L}_{l,i}\hat{\rho}\otimes|0\rangle\langle 0|\Bigg{\}}.$ (63)
Let us now evaluate the second term in Eq. (B.2). To illustrate the derivation in a concise and transparent manner, we restrict our attention to the contribution in ±$\mathcal{L}_{\pm,i}$ that is proportional to the first derivative of the drive; the remaining terms can be treated analogously. Focusing on this component, one readily finds
±$\mathcal{L}_{\pm,i}\hat{\rho}\otimes|0\rangle\langle 0|=-i\eta_{i}\Big{(}$ $\Omega_{i}^{\prime}\hat{b}_{i}^{\pm}\hat{\rho}\otimes\hat{\sigma}_{i}^{\dagger}|0\rangle\langle 0|$
@@
where we have defined the coefficients $c_{\lambda,ij}=\phi_{\lambda,i}\phi_{\lambda,j}/\sum_{m}\phi_{\lambda,m}^{2}$. While the spin-motion interaction induced by the driving field is local (i. e., it acts individually on each emitter), the delocalized character of the spin eigenstates introduces non-locality into the dynamics. This effect is evidenced by the appearance of an additional atomic index $j$, and ultimately gives rise to non-local interactions between the motional degrees of freedom.
@@
[…]
Finally, we apply the superoperator $\mathcal{L}_{-l,j}$ to the result of Eq. (B.2) after the integration in Eq. (67), and trace over the spin degrees of freedom. This yields non-zero contributions only from terms of the form $\mathrm{Tr}_{S}\{|0\rangle\langle 0|\}$ and $\mathrm{Tr}_{S}\{\hat{\sigma}_{i}^{\dagger}|0\rangle\langle 0|\hat{\sigma}_{i}\}$. Upon additionally performing the rotating wave approximation (RWA) on the motional Liouvillian $\mathcal{L}_{M}(t)\hat{\rho}$ in the interaction picture, Eq. (B.2) reduces to the final master equation governing the atomic motion given in Eq. (16).
@@ C.1 Lamb-Dicke approximation
To avoid an heavy notation, we are using the same labels for the different contribution in Eq. (73) as the one we used in the main text, although the terms here include motion along different directions.
@@
where $\nu_{j\alpha}$ is the trap frequency at site $j$ along the spatial direction $\alpha\in\{x,y,z\}$. We defined the bosonic annihilation operator $\hat{b}_{j\alpha}$ for a motional quantum at site $j$ along the direction $\alpha$, as well as the corresponding dimensionless position operators $\hat{R}_{j\alpha}\equiv\sqrt{\bar{\nu}/\nu_{j\alpha}}(\hat{b}^{\dagger}_{j\alpha}+\hat{b}_{j\alpha})$. Here, the average trap frequency is defined as $\bar{\nu}\equiv\sum_{j}\sum_{\alpha}\nu_{j\alpha}/3N$, from which we introduce the averaged Lamb-Dicke parameter $\eta\equiv\sqrt{\nu_{R}/\bar{\nu}}$. We further define the site- and direction-dependent Lamb-Dicke parameter as $\eta_{j\alpha}=\sqrt{\nu_{R}/\nu_{j\alpha}}$.
@@ C.2 Effective master equation for the atomic motion
$\begin{split}\frac{d}{dt}\hat{\rho}=&-i\Big{[}\sum_{i,j}\sum_{\alpha,\beta}V_{i\alpha j\beta}e^{\mathrm{i}(\nu_{i\alpha}-\nu_{j\beta})t}\hat{b}^{\dagger}_{i\alpha}\hat{b}_{j\beta},\hat{\rho}\Big{]}\\ +&\sum_{i,j}\sum_{\alpha,\beta}\!\mathcal{R}_{i\alpha j\beta}^{-}e^{\mathrm{i}(\nu_{i\alpha}-\nu_{j\beta})t}\Big{(}\hat{b}_{j\beta}\hat{\rho}\hat{b}_{i\alpha}^{\dagger}\!-\!\frac{1}{2}\{\hat{b}_{i\alpha}^{\dagger}\hat{b}_{j\beta},\hat{\rho}\}\Big{)}\\ +&\sum_{i,j}\sum_{\alpha,\beta}\!\mathcal{R}_{i\alpha j\beta}^{+}e^{-\mathrm{i}(\nu_{i\alpha}-\nu_{j\beta})t}\Big{(}\hat{b}_{j\beta}^{\dagger}\hat{\rho}\hat{b}_{i\alpha}\!-\!\frac{1}{2}\{\hat{b}_{i\alpha}\hat{b}_{j\beta}^{\dagger},\hat{\rho}\}\Big{)}.\end{split}$ (80)
@@ Appendix D Numerical simulation of the cooling dynamics
In Fig. 8, we compare the steady-state phonon population for a chain of $N$ atoms obtained from two approaches: (i) solving the rate equations derived from the effective master equation Eq. (16), and (ii) numerically simulating the full Lamb-Dicke master equation Eq. (7) using the method described in this section. The excellent agreement between the two methods validates the accuracy of the effective model.
Figure 8: Steady state population $\bar{n}$ for a 1D array of up to $N=7$ atoms moving perpendicular to the array direction. Different colors represent different lattice spacing as specified by the legend. Top and bottom panel refer to the case where $\bar{\nu}=20\gamma_{0}$ and $\bar{\nu}=0.25\gamma_{0}$ respectively. We consider $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ in both panels, such that the system is in the array cooling regime. The empty circles represent the result obtain from the effective master equation Eq. (16), while filed diamond are results from the numerical simulation of the Lamb-Dicke master equataion Eq. (7). Due to the large steady-state phonon populations for $d=0.3\lambda_{0}$ and $d=0.4\lambda_{0}$ at $\bar{\nu}=0.25\gamma_{0}$, numerical simulations of Eq. (7) require cutoffs $N_{\text{cutoff}}$ beyond our computational capabilities and were therefore not performed.
@@ E.1 Dependence on the transition dipole moment
Figure 9: Dependence on atomic polarization. (a) Sideband structure for atoms polarized perpendicular to the axis of the chain. The antisymmetric cooling sideband, symmetric heating sidebands, and the recoil Lorentzian cannot overlap as $d$ decreases. As a result, the phonon number $\bar{n}$ remains below that of non-interacting emitters for all $\bar{\nu}$, both for (b) chains of 2 atoms and (c) chains of 4 atoms. (d) Sideband structure for atoms polarized parallel to the axis of the chain. The antisymmetric cooling sideband can become resonant with symmetric heating sidebands and the recoil Lorentzian. The steady state phonon number $\bar{n}$ for (e) chains of 2 atoms and (f) chains of 4 atoms can become larger than that of non-interacting atoms for specific combinations of $\bar{\nu}$ and $J_{12}$. The solid lines correspond to a drive on resonance with the cooling sideband of the collective spin mode that leads to a smallest phonon number, whereas the dashed lines correspond to a drive on resonance with the cooling sideband of the darkest spin mode. Notably, when the cooling sideband of the darkest spin mode becomes resonant with either a heating sideband of another spin mode or the recoil Lorentzian, cooling can be enhanced by driving a spin mode with a larger linewidth for which this coalescence does not occur. The results are obtained for $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=5\times 10^{-4}\gamma_{0}$.
@@
In this section, we discuss the dependence of the steady state phonon number attained in the array cooling regime on the direction of atomic polarization. For that, it is instructive to first consider the case of two atoms located at a distance $d$ along the $x$-axis and moving along the $z$-direction. The optical response of the system features: (a) two cooling sidebands associated to the antisymmetric and symmetric collective spin modes, respectively centered at $\Delta=-\nu_{1}-J_{12}$ and $\Delta=-\nu_{1}+J_{12}$; (b) two heating sidebands associated to the antisymmetric and symmetric collective spin modes, respectively centered at $\Delta=\nu_{1}-J_{12}$ and $\Delta=\nu_{1}+J_{12}$; (c) a heating contribution from the recoil or diffusive motion in the form a Lorentzian centered at $\Delta=J_{12}$. As we have done throughout this work, we consider the drive to be on resonance with the antisymmetric cooling sideband. Then, the symmetric heating sideband and the contribution from recoil are respectively detuned from the driving field by $\delta_{hs}=2(\nu_{1}+J_{12})$ and $\delta_{r}=\nu_{1}+2J_{12}$. There exist two different scenarios depending on the sign of the coherent dipole-dipole interaction $J_{12}$ at small inter-atomic distances, which in turn depends on the atomic polarization [Fig. 9(a) and (d)].
@@
For polarization perpendicular to the array and the drive (i. e., along $\mathbf{y})$, we obtain $J_{12}>0$ for $d<0.7\lambda_{0}$. Since both the trap frequency $\nu_{j}$ and $J_{12}$ are positive, the detunings $\delta_{hs}$ and $\delta_{r}$ cannot become zero. In other words, the symmetric heating sideband and the heating contribution arising from recoil can never be on resonance with the antisymmetric cooling sideband, as illustrated in Fig. 9(a). As a result, the steady state photon number in the array cooling regime and at small distances $d$ is always smaller than that of non-interacting atoms, as shown in Fig. 9(b). Note that this is the configuration considered in the main text.
@@
For atoms polarized along the direction of the atomic chain (i. e., along $\mathbf{x})$, we obtain $J_{12}<0$ for $d<0.43\lambda_{0}$. As illustrated in Fig. 9(d), it is now possible for the antisymmetric cooling sideband to become resonant with either the symmetric heating sideband or the recoil Lorentzian. This respectively occurs when $\nu_{1}=-J_{12}$ and $\nu_{1}=-2J_{12}$, and leads to increased steady state phonon numbers as shown in Fig. 9(e). In particular, the phonon number can become larger than that of non-interacting atoms, such that the cooling advantage arising from the narrow collective resonance is lost. In general, this effect can be minimized by driving on resonance with the cooling sideband associated to another spin mode (in this case, the symmetric cooling sideband), or by separating the atoms by distances such that the different resonances do not overlap for the specific trap frequency considered.
@@
This phenomenon also occurs for larger atom numbers. In Fig. 9(c) and (f), we show an example for a chain of $4$ atoms. For perpendicular polarization, the most subradiant spin mode is also associated to the most negative energy shift, and efficient cooling is attained for all spacings and trap frequencies. For polarization along the atomic chain, that is no longer the case and cooling via the most subradiant spin mode leads to phonon numbers larger than that of non-interacting atoms for certain combinations of $\bar{\nu}$ and $d$ that lead to coalescence of sidebands. Again, it is generally possible to minimize this extra heating by driving on resonance with the cooling sideband associated to another collective spin mode for which this coalescence does not occur [see the reduced temperature for the solid lines as opposed to the dashed lines in Fig. 9(f)]. Alternatively, there always exists a spacing $d$ such that the array can be cooled to temperatures lower than that of non-interacting atoms independently of $\bar{\nu}$.

### v3

@@ Acknowledgements.
O.R.-B. acknowledges support from Fundación Mauricio y Carlota Botton and from Fundació Bancaria “la Caixa” (LCF/BQ/AA18/11680093). SFY thanks the NSF throughPHY-2207972, the CUA PFC PHY-2317134, and the Q-SEnSE QLCI OMA-2016244. A. A.-G. acknowledges support from the AFOSR through their Young Investigator Prize (No. 21RT0751), from the NSF through their CAREER Award (No. 2047380), and from the David and Lucile Packard Foundation. C.C.R. acknowledges support from the European Union’s Horizon Europe program under the Marie Sklodowska Curie Action LIME (Grant No. 101105916). This research was supported in part by the Austrian Science Fund (FWF) [10.55776/COE1 QuantA] and by the grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
@@ Appendix A Free-space dipole-dipole interactions
[…]
@@ Appendix B Derivation of the effective master equation
In this appendix, we derive the effective master equation (16) describing the laser cooling of the motional degrees of freedom of the system. The derivation relies on two approximations. First, we assume the internal or spin degrees of freedom to be only weakly excited, which is valid for weak and off-resonant drives. This allows to approximate the optical response of the atomic ensemble by its linear response. Second, that the coupling between internal and external degrees of freedom is smaller than the characteristic timescale of the spin dynamics, which allows for an adiabatic elimination of the spins. Below, we derive specific condition for the validity of these two assumptions. Note that the derived effective master equation includes the effects arising from light-induced dipole-dipole interactions mediated by the vacuum electromagnetic field, and therefore describes collective cooling in dense atomic ensembles.
@@ B.1 Displacement transformation in the low spin-excitation regime
The starting point of our analysis is the master equation for the full system in the Lamb-Dicke approximation, Eq. (7). Under the assumption of weak saturation, $\langle\hat{\sigma}_{i}^{\dagger}\hat{\sigma}_{i}\rangle\ll 1$ $\forall i$, we approximate the spin operators by bosonic operators with $\hat{\sigma}_{i},\hat{\sigma}_{j}^{\dagger}]=\delta_{ij}$. Next, we apply the unitary transformation
$\hat{U}_{d}=\exp\left(\sum_{j}(s_{j}^{*}\hat{\sigma}_{j}+\beta_{j}^{*}\hat{b}_{j}-\text{H.c.})\right),$ (47)
which displaces the bosonic spin and phonon operators as $\hat{\sigma}_{i}\rightarrow\hat{\sigma}_{i}+s_{i}$ and $\hat{b}_{i}\rightarrow\hat{b}_{i}+\beta_{i}$. Equation (47) corresponds to a shift of system’s degrees of freedom to a new equilibrium position determined by the strength of the external driving laser. The displacements can be approximated as a power series in the Lamb-Dicke parameter. To leading order they read
@@
where $\hat{H}_{S}$ is defined in Eq. (10) and the last term represents and additional correction of order $\eta^{2}$ arising from the displacement of the atomic position. Note that even for $\beta_{i}^{(1)}=0$, $\hat{H}_{S}$ contains a second order corrections in $\eta$ that arises from the zero-point motion of the emitters. Formally, this correction is obtained from terms of the form $\hat{b}_{i}\hat{b}_{i}^{\dagger}$, which become $\hat{b}_{i}^{\dagger}\hat{b}_{i}+1$ upon normal ordering (see discussion at the end of Sec. II.1). Finally, the population recycling term reads
@@
where $\mathcal{L}_{\pm}\equiv-(i/\hbar)\left[\hat{H}_{\pm},\hat{\mu}\right]+\mathcal{D}_{\pm}\hat{\mu}$. To first order in $\eta$, the interaction reads
@@
[…]
@@ B.2 Adiabatic elimination of spin degrees of freedom
$\eta_{i}\Omega_{i},\eta_{i}J_{ij}^{\prime}|s_{j}^{(0)}|,\eta_{i}\gamma_{ij}^{\prime}|s_{j}^{(0)}|\ll\nu_{i},\gamma_{\lambda}\quad\forall i,j,\lambda.$ (60)
@@
where we defined $\mathcal{L}_{\pm}=\sum_{i}\mathcal{L}_{\pm,i}$, with $\mathcal{L}_{\pm,i}$ containing only motional operators of emitter $i$.
[…]
$\frac{d}{dt}\mathcal{P}\hat{\mu}=\mathcal{L}_{M}(t)\mathcal{P}\hat{\mu}+\mathcal{L}_{\mathrm{int}}(t)\!\!\int_{0}^{\infty}\!\!\!d\tau\,e^{\tau\mathcal{L}_{S}}\mathcal{L}_{\mathrm{int}}(t-\tau)\mathcal{P}\hat{\mu}$. (62)
@@
$\times\int_{0}^{\infty}d\tau\,e^{\tau(\mathcal{L}_{S}-il\nu_{i})}\mathcal{L}_{l,i}\hat{\rho}\otimes|0\rangle\langle 0|\Bigg{\}}.$ (63)
Let us now evaluate the second term in Eq. (B.2). To illustrate the derivation in a concise and transparent manner, we restrict our attention to the contribution in $\mathcal{L}_{\pm,i}$ that is proportional to the first derivative of the drive; the remaining terms can be treated analogously. Focusing on this component, one readily finds
$\mathcal{L}_{\pm,i}\hat{\rho}\otimes|0\rangle\langle 0|=-i\eta_{i}\Big{(}$ $\Omega_{i}^{\prime}\hat{b}_{i}^{\pm}\hat{\rho}\otimes\hat{\sigma}_{i}^{\dagger}|0\rangle\langle 0|$
@@
where we have defined the coefficients $c_{\lambda,ij}=\phi_{\lambda,i}\phi_{\lambda,j}/\sum_{m}\phi_{\lambda,m}^{2}$. While the spin-motion interaction induced by the driving field is local (i. e., it acts individually on each emitter), the delocalized character of the spin eigenstates introduces non-locality into the dynamics. This effect is evinced by the appearance of an additional atomic index $j$, and ultimately gives rise to nonlocal interactions between the motional degrees of freedom.
@@
$\int_{0}^{\infty}d\tau\,e^{-i\tau(\epsilon_{\lambda}\pm\nu_{i})}=\frac{-i}{\epsilon_{\lambda}\pm\nu_{i}}\equiv E_{\lambda i\pm}.$ (67)
Finally, we apply the superoperator $\mathcal{L}_{-l,j}$ to the result of Eq. (B.2) after the integration in Eq. (67), and trace over the spin degrees of freedom. This yields nonzero contributions only from terms of the form $\mathrm{Tr}_{S}\{|0\rangle\langle 0|\}$ and $\mathrm{Tr}_{S}\{\hat{\sigma}_{i}^{\dagger}|0\rangle\langle 0|\hat{\sigma}_{i}\}$. Upon additionally performing the rotating wave approximation (RWA) on the motional Liouvillian $\mathcal{L}_{M}(t)\hat{\rho}$ in the interaction picture, Eq. (B.2) reduces to the final master equation governing the atomic motion given in Eq. (16).
@@ C.1 Lamb-Dicke approximation
To avoid heavy notation, we are using the same labels for the different contribution in Eq. (73) as the one we used in the main text, although the terms here include motion along different directions.
@@
where $\nu_{j\alpha}$ is the trap frequency at site $j$ along the spatial direction $\alpha\in\{x,y,z\}$. We defined the bosonic annihilation operator $\hat{b}_{j\alpha}$ for a motional quantum at site $j$ along direction $\alpha$, as well as the corresponding dimensionless position operators $\hat{R}_{j\alpha}\equiv\sqrt{\bar{\nu}/\nu_{j\alpha}}(\hat{b}^{\dagger}_{j\alpha}+\hat{b}_{j\alpha})$. Here, the average trap frequency is defined as $\bar{\nu}\equiv\sum_{j}\sum_{\alpha}\nu_{j\alpha}/3N$, from which we introduce the averaged Lamb-Dicke parameter $\eta\equiv\sqrt{\nu_{R}/\bar{\nu}}$. We further define the site- and direction-dependent Lamb-Dicke parameter as $\eta_{j\alpha}=\sqrt{\nu_{R}/\nu_{j\alpha}}$.
@@ C.2 Effective master equation for the atomic motion
[…]
@@ Appendix D Numerical simulation of the cooling dynamics
In Fig. 8, we compare the steady-state phonon population for a chain of $N$ atoms obtained from two approaches: (i) solving the rate equations derived from the effective master equation (16), and (ii) numerically simulating the full Lamb-Dicke master equation (7) using the method described in this section. The excellent agreement between the two methods validates the accuracy of the effective model.
Figure 8: Steady-state population $\bar{n}$ for a one-dimensional array of up to $N=7$ atoms moving perpendicular to the array direction. Different colors represent different lattice spacing as specified by the legend. Top and bottom panel refer to the case where $\bar{\nu}=20\gamma_{0}$ and $\bar{\nu}=0.25\gamma_{0}$ respectively. We consider $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ in both panels, such that the system is in the array cooling regime. The empty circles represent the result obtain from the effective master equation (16), while filled diamond are results from the numerical simulation of the Lamb-Dicke master equataion (7). Due to the large steady-state phonon populations for $d=0.3\lambda_{0}$ and $d=0.4\lambda_{0}$ at $\bar{\nu}=0.25\gamma_{0}$, numerical simulations of Eq. (7) require cutoffs $N_{\text{cutoff}}$ beyond our computational capabilities and were therefore not performed.
@@ E.1 Dependence on the transition dipole moment
Figure 9: Dependence on atomic polarization. (a) Sideband structure for atoms polarized perpendicular to the axis of the chain. The antisymmetric cooling sideband, symmetric heating sidebands, and the recoil Lorentzian cannot overlap as $d$ decreases. As a result, the phonon number $\bar{n}$ remains below that of noninteracting emitters for all $\bar{\nu}$, both for (b) chains of two atoms and (c) chains of four atoms. (d) Sideband structure for atoms polarized parallel to the axis of the chain. The antisymmetric cooling sideband can become resonant with symmetric heating sidebands and the recoil Lorentzian. The steady state phonon number $\bar{n}$ for (e) chains of two atoms and (f) chains of four atoms can become larger than that of noninteracting atoms for specific combinations of $\bar{\nu}$ and $J_{12}$. The solid lines correspond to a drive on resonance with the cooling sideband of the collective spin mode that leads to a smallest phonon number, whereas the dashed lines correspond to a drive on resonance with the cooling sideband of the darkest spin mode. Notably, when the cooling sideband of the darkest spin mode becomes resonant with either a heating sideband of another spin mode or the recoil Lorentzian, cooling can be enhanced by driving a spin mode with a larger linewidth for which this coalescence does not occur. The results are obtained for $\eta=0.02$ and $\nu_{j}=\bar{\nu}+\delta\nu(j-\lfloor N/2\rfloor-1)$ with $\bar{\nu}=20\gamma_{0}$ and $\delta\nu=5\times 10^{-4}\gamma_{0}$.
@@
In this section, we discuss the dependence of the steady state phonon number attained in the array cooling regime on the direction of atomic polarization. For that, it is instructive to first consider the case of two atoms located at a distance $d$ along the $x$-axis and moving along the $z$-direction. The optical response of the system features: (a) two cooling sidebands associated to the antisymmetric and symmetric collective spin modes, respectively centered at $\Delta=-\nu_{1}-J_{12}$ and $\Delta=-\nu_{1}+J_{12}$; (b) two heating sidebands associated to the antisymmetric and symmetric collective spin modes, respectively centered at $\Delta=\nu_{1}-J_{12}$ and $\Delta=\nu_{1}+J_{12}$; (c) a heating contribution from the recoil or diffusive motion in the form a Lorentzian centered at $\Delta=J_{12}$. As we have done throughout this work, we consider the drive to be on resonance with the antisymmetric cooling sideband. Then, the symmetric heating sideband and the contribution from recoil are respectively detuned from the driving field by $\delta_{hs}=2(\nu_{1}+J_{12})$ and $\delta_{r}=\nu_{1}+2J_{12}$. There exist two different scenarios depending on the sign of the coherent dipole-dipole interaction $J_{12}$ at small interatomic distances, which in turn depends on the atomic polarization [Fig. 9(a) and (d)].
@@
For polarization perpendicular to the array and the drive (i. e., along $\mathbf{y})$, we obtain $J_{12}>0$ for $d<0.7\lambda_{0}$. Since both the trap frequency $\nu_{j}$ and $J_{12}$ are positive, the detunings $\delta_{hs}$ and $\delta_{r}$ cannot become zero. In other words, the symmetric heating sideband and the heating contribution arising from recoil can never be on resonance with the antisymmetric cooling sideband, as illustrated in Fig. 9(a). As a result, the steady state photon number in the array cooling regime and at small distances $d$ is always smaller than that of noninteracting atoms, as shown in Fig. 9(b). Note that this is the configuration considered in the main text.
@@
For atoms polarized along the direction of the atomic chain (i. e., along $\mathbf{x})$, we obtain $J_{12}<0$ for $d<0.43\lambda_{0}$. As illustrated in Fig. 9(d), it is now possible for the antisymmetric cooling sideband to become resonant with either the symmetric heating sideband or the recoil Lorentzian. This respectively occurs when $\nu_{1}=-J_{12}$ and $\nu_{1}=-2J_{12}$ and leads to increased steady-state phonon numbers as shown in Fig. 9(e). In particular, the phonon number can become larger than that of noninteracting atoms, such that the cooling advantage arising from the narrow collective resonance is lost. In general, this effect can be minimized by driving on resonance with the cooling sideband associated to another spin mode (in this case, the symmetric cooling sideband), or by separating the atoms by distances such that the different resonances do not overlap for the specific trap frequency considered.
@@
This phenomenon also occurs for larger atom numbers. In Fig. 9(c) and (f), we show an example for a chain of four atoms. For perpendicular polarization, the most subradiant spin mode is also associated to the most negative energy shift, and efficient cooling is attained for all spacings and trap frequencies. For polarization along the atomic chain, that is no longer the case and cooling via the most subradiant spin mode leads to phonon numbers larger than that of noninteracting atoms for certain combinations of $\bar{\nu}$ and $d$ that lead to coalescence of sidebands. Again, it is generally possible to minimize this extra heating by driving on resonance with the cooling sideband associated to another collective spin mode for which this coalescence does not occur [see the reduced temperature for the solid lines as opposed to the dashed lines in Fig. 9(f)]. Alternatively, there always exists a spacing $d$ such that the array can be cooled to temperatures lower than that of noninteracting atoms independently of $\bar{\nu}$.